# Macroeconomic Spillovers of Weather Shocks across U.S. States

### v2

@@
Abstract: We estimate the short-run effects of severe weather shocks on local economic activity and cross-border spillovers operating through economic linkages between U.S. states. We measure weather shocks using emergency declarations triggered by natural disasters and estimate their impacts with a monthly Global Vector Autoregressive (GVAR) model for U.S. states. Impulse responses highlight country-wide effects of weather shocks hitting individual regions. Taking into account economic interconnections between states allows capturing much stronger spillovers than those associated with mere spatial adjacency. Also, geographical heterogeneity is critical for assessing country-wide effects of weather shocks, and network effects amplify the local impacts of shocks.
@@
JEL Codes: C32, R11, Q51, Q54. (∗) Corresponding author: Graziano Moramarco, Department of Economics, University of Bologna, Piazza Scaravilli 2, Bologna, Italy. Email: graziano.moramarco@unibo.it. Acknowledgments: For this research, Graziano Moramarco gratefully acknowledges funding from the European Union and Ministero dell’Università e della Ricerca (FSE REACT EU - PON Ricerca e Innovazione 2014-2020).
@@ 1 Introduction
Climate-related risks are now regarded as major sources of economic vulnerability. These risks are associated with adverse climatic events (physical risks) or induced by the process of adjustment towards a low-carbon economy (transition risks). Extreme weather events are the primary types of physical risks and are becoming more frequent as a result of climate change (IPCC, 2023). Evaluating their potentially long-lasting economic impacts (Batten, , 2018; Hsiang and Jina, , 2014) requires considering not only local effects but also spillovers across interconnected economic systems (Benzie et al., , 2021; Cashin et al., , 2017).
In this study we examine the impact of severe weather shocks on state-level economic activity in the United States, considering the cross-border spillovers resulting from economic interconnections among states and regions. We rely on a monthly Global Vector Autoregressive (GVAR) model (Pesaran et al., , 2004) that captures economic interrelationships between states and regions using bilateral trade flows within the U.S. and include a proxy of severe weather shocks, based on a detailed database of emergency declarations triggered by natural disasters.
@@
We find that weather shocks generate direct negative macroeconomic effects at the local level and significant negative spillovers to U.S. regions not directly affected. Furthermore, we perform several robustness checks to evaluate the quantitative and qualitative relevance of our results. We first assess the importance of considering economic interconnections rather than mere spatial proximity of states, when measuring spillovers. Therefore, we compare our baseline GVAR model with an alternative GVAR specification where states are connected by a spatial adjacency matrix. Moreover, our model allows for parameter heterogeneity between states. We thus evaluate the importance of this feature by comparing our GVAR model with its “homogeneous-parameter counterpart”, namely a spatial panel dynamic model. Next, we consider a country-wide regression for the U.S. to check whether disaggregation at the state level matters at all. We show that disaggregation, parameter heterogeneity and economic linkages between states are key for quantifying spillover effects and the overall impact of weather shocks in the U.S. Lastly, we provide a measure of the impact of spillovers for each state, by comparing responses to weather shocks between the GVAR model – that allows for spillovers – and individual state-specific Autoregressive Distributed Lag (ARDL) models where spillovers are implicitly muted.
@@
The paper is related to several streams of literature. First, it adds to the rapidly growing literature on the macroeconomic effects of extreme weather events and climate change (see e.g. Boldin and Wright, , 2015; Burke et al., , 2015; Colacito et al., , 2019; Dell et al., , 2014; Colombo and Ferrara, , 2024; Billio et al., , 2020; Ahmadi et al., , 2022). In particular, it contributes to the strand of research on cross-border impacts of climate change (Benzie et al., , 2021; Carter et al., , 2021; Feng et al., , 2023). There is ample evidence that the effects of natural disasters propagate through international trade and production networks (Barrot and Sauvagnat, , 2016; Boehm et al., , 2019; Carvalho et al., , 2020; Feng et al., , 2023; Forslid and Sanctuary, , 2023; Kashiwagi et al., , 2021). Our paper is also related to the literature on climate econometrics (see e.g. Burke et al., , 2015; Hsiang, , 2016; Kahn et al., , 2021). Focusing on the methodology, it is worth pointing out that the GVAR framework has been used to investigate the international spillovers of El Niño weather shocks by Cashin et al., (2017).
@@
Five papers, to the best of our knowledge, are closely related to our work, in that they have focused on the effects of climate and weather extremes on the U.S. economy. Mohaddes et al., (2022) and Natoli, (2023) investigate the macroeconomic effects of climate change across the U.S. over the periods 1963–2016 and 1970-2019, respectively. Unlike our analysis, these authors focus on temperature and precipitation anomalies rather than natural disasters and do not assess spillover effects between states. Kim et al., (2021) investigate the effects of severe weather conditions, proxied by the Actuaries Climate Index, on the aggregate U.S. economy over the period 1963-2019. They find that the effects have become significant only in recent years. Eickmeier et al., (2024) analyze the impact of natural disasters on the aggregate U.S. economy and financial markets using monthly data from 2000 to 2019, and focusing on extreme temperature events, floods, and storms. Tran and Wilson, (2020) study the long-run impact of natural disasters on income using a similar proxy based on U.S. disaster declarations. Our paper relies on monthly data and focuses on short-run impacts, thus we complement their analysis based on annual data and concentrating on long-run impacts. Moreover, we study spillover effects based on economic interdependence, while Tran and Wilson, (2020) focus on spatial spillovers.
@@
The rest of the paper is organized as follows. Section 2 presents the data, with a particular focus on our proxy for severe weather shocks, Section 3 discusses the econometric model, Section 4 presents the results, and Section 5 concludes. A Supplement with further results completes the paper.
@@ 2 Data
2.1 Weather shocks
We build a novel monthly proxy of weather shocks using the Disaster Declarations Summary dataset maintained by the U.S. Federal Emergency Management Agency (FEMA). FEMA, established in 1979 with an executive order signed by President Carter, is part of the Department of Homeland Security and is responsible for coordinating the federal government’s relief efforts after natural or man-made domestic disasters. All federally declared disasters since 1953 are collected in the Disaster Declarations Summary. For each declaration, the database provides detailed information on the states and counties affected by the disaster, its beginning and end dates, as well as on the type of disaster.
@@ 2.1 Weather shocks
Figure 1: Number of emergency declarations by type and season: Jan. 1990 to Dec. 2019
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Notes: the maps show the count of emergency declarations (total and by type of event). Colors range from lighter to darker as the count of events grows. Hawaii and Alaska are omitted for better display.
We focus on declarations caused by events that are directly related to weather conditions. In particular, we consider the following main types of events that group the weather-related declarations in FEMA’s database: winter weather (i.e., freezing, ice storms, snow), tropical storms (i.e., hurricanes, typhoons and tornadoes), storms (i.e., coastal and severe storms), fire, flood, drought. Figure 1 displays the events examined in our analysis. The left panel shows that storms and fires are associated with the majority of declarations. In the right panel of the figure, we report the distribution of events over different seasons that highlights that most events exhibit the expected seasonal patterns. Figure 2 shows the count of declarations by state and by type. Darker colors indicate a higher number of events. If we do not distinguish by event type – upper left panel – we can see that emergency declarations are evenly distributed across states. The same holds true for storms and floods. On the contrary, declarations triggered by severe winter weather, tropical storms, fires, and drought are geographically concentrated in the northeast, southeast, and west of the U.S., respectively. Alternative representations of the maps in Figure 2, based either on the count of declarations weighted by the state population density or aggregated over U.S. climate regions, are provided in the online Appendix.
Measuring weather shocks with disaster declarations. We construct our proxy of weather shocks as follows. First, for each state $i$ we create a dummy variable that takes value one if a weather-related disaster was declared in month $t$, and zero otherwise ($emergency_{it}$). Second, we multiply this dummy variable by a weight, given by the ratio between the number of counties affected by the disaster ($hit_{it}$) and the number of counties (or county equivalents) in state $i$ ($counties_{i}$):
@@
Therefore, our state-level weather shock has unit value only if in a given month emergency declarations are issued for all the counties in a state. To calculate an aggregate weather shock for the U.S., we consider the number of counties hit as a share of the total number of counties in the 50 states considered in the analysis ($N_{c}=3142$): [^2] The list of U.S. counties and county equivalents is sourced from: https://en.wikipedia.org/wiki/County.
@@
The weather shock series for the U.S. appears in Figure 3 where we highlight the five largest shocks over the period 1990-2019. Hurricane Katrina is by far the largest event, as measured by the share of counties where emergency declarations have been issued. Figure 3 also shows that the number of declarations for the U.S. is highly volatile, but there is no visible trend, nor structural breaks in the series.
Figure 3: Weather shocks for the US
@@
Notes: for Hurricane Katrina, emergency declarations were issued in almost all states. In states not directly hit by the hurricane the declarations go under the header “Hurricane Katrina Evacuation”.
We have considered different weighting schemes that aim to allocate more weight to areas with higher economic activity. These alternative weights are based either on resident population by state from the 2010 U.S. Census or on state-level employment rates from the Bureau of Labor Statistics. See also Boldin and Wright, (2015); Colacito et al., (2019). Results based on these proxies are qualitatively similar to those obtained relying on the share of counties hit by a natural disaster. This does not come as a surprise, given that the GVAR captures economic linkages between states and hence implicitly assigns more weight to events in hitting states that are central to the U.S. economy. [^3] Results are available from the authors upon request.
Comparison with other proxies of weather shocks. Other variables can be used to capture the effects of severe weather conditions on macroeconomic aggregates. The first option is represented by the use of weather data, which is common in the literature (see e.g. Bloesch and Gourio, , 2015; Boldin and Wright, , 2015; Colacito et al., , 2019; Wilson, , 2016; van der List and Wilson, , 2016). It is well known that measuring the weather conditions raises several problems, such as the heterogeneous quality, availability and representativeness of data from different weather stations (Dell et al., , 2014). Even after measurement issues have been taken care of, it is not clear which variables best capture weather shocks. Severe winter weather is a case in point: snowfall, temperatures, wind speed and heating degree days are just an incomplete list of alternatives available to researchers. Moreover, the societal impacts of weather events depend on the degree of preparedness of the area hit by the event. Therefore, relying on weather variables is likely to introduce some measurement error and omitted variable bias in the empirical analysis, and does not guarantee to correctly identify weather events with sizable impacts on the U.S. economy.
An alternative that goes some way towards improving the identification of such events is to rely on databases explicitly designed to capture the societal impacts of weather, such as the Billion–Dollar Weather and Climate Disasters dataset. This dataset, maintained by the National Oceanic and Atmospheric Administration (NOAA, 2024), is tightly linked to our proxy. [^4] Datasets of this type are often used in empirical analyses focusing on the economic impacts of natural disasters (Barro, , 2006; Gourio, , 2012). Recent empirical studies relying on this kind of data include Baker and Bloom, (2013), Cavallo et al., (2013), Strömberg, (2007) who rely on the database maintained by Centre for Research on the Epidemiology of Disasters at the University of Louvain, Belgium. As its name suggests, the Billion–Dollar Weather and Climate Disasters dataset arbitrarily focuses only on disasters with an overall cost exceeding the threshold of one billion dollar and hence leads to a series with very low variability that might miss smaller, yet economically important events.
Yet another option, used by Boldin and Wright, (2015), is to focus on data about weather-affected jobs (i.e., the time series of absences from work due to bad weather), collected by the Bureau of Labor Statistics in the Current Population Survey. This series has three shortcomings: it does not allow us to isolate work absences due to severe winter weather from absences due to hurricanes or other events. Moreover, it only captures temporary effects of weather events and might miss economically relevant episodes that do not cause work absences but impact the supply chain. Last but not least, this measure is not suited for measuring spillovers across states: in fact absences from work in one region do not affect presence at work in other regions.
Figure 4 compares our proxy with the number of weather-affected jobs and the count of Billion Dollar Disasters. We carried out a sample cross-correlation analysis to study the lead-lag relationship between different proxies: the contemporaneous sample correlation between our proxy, the count of weather-affected jobs, and the number of Billion Dollar Disasters is equal to 0.14 and 0.01 (this becomes 0.11 if we consider the log-count), respectively. In both cases, there is no clear indication of a lead-lag relationship between the two series.
Figure 4: Alternative proxies of weather shocks for the US
Refer to captionRefer to caption
@@
To sum up, the FEMA dataset includes all disasters “of such severity and magnitude that effective response is beyond the capabilities of the state ()($\ldots$)( … ) and that supplemental federal assistance is necessary.” [^5] Source: https://www.fema.gov/disaster/how-declared Extreme weather events measured in this way represent “proper” shocks in the parlance of Ramey, (2016): (i) they are exogenous with respect to the current and lagged endogenous variables of the model; (ii) they are uncorrelated with other exogenous shocks (e.g., technological, monetary or fiscal policy shocks), and (iii) they are generally unanticipated.

### v3

@@
Abstract: We estimate the short-run effects of weather-related disasters on local economic activity and cross-border spillovers that operate through economic linkages between U.S. states. To this end, we use emergency declarations triggered by natural disasters and estimate their effects using a monthly Global Vector Autoregressive (GVAR) model for U.S. states. Impulse responses highlight the nationwide effects of weather-related disasters that hit individual regions. Taking into account economic linkages between states allows capturing much stronger spillovers than those associated with mere spatial proximity. The results underscore the importance of geographic heterogeneity for impact evaluation and the critical role of supply-side propagation mechanisms.
@@
JEL Codes: C32, E32, R11, Q54. (∗)Corresponding author: Graziano Moramarco, Department of Economics, University of Bologna, Piazza Scaravilli 2, Bologna, Italy. Email: graziano.moramarco@unibo.it. Acknowledgments: Emanuele Bacchiocchi gratefully acknowledges financial support from Ministero dell’Università e della Ricerca (MUR), PRIN 20229PFAX5, and the University of Bologna, RFO grants. Andrea Bastianin acknowledges financial support within the “Fund for Departments of Excellence academic funding” provided by MUR, established by Stability Law, namely “Legge di Stabilita n.232/2016, 2017” - Project of the Department of Economics, Management, and Quantitative Methods, University of Milan. Graziano Moramarco gratefully acknowledges funding from the European Union and MUR (REACT-EU PON Ricerca e Innovazione 2014-2020 D.M. 1062/2021 CUP J41B21012140007 and GRINS project, code PE0000018 CUP J33C22002910001).
@@ 1 Introduction
Climate-related risks are now regarded as major sources of economic vulnerability. These risks are associated with adverse climatic events (physical risks) or induced by the process of adjustment towards a low-carbon economy (transition risks). Extreme weather events are the main type of physical risk, and at least some of these events have increased in frequency and severity in the U.S. [^1] The Fifth National Climate Assessment (USGCRP, , 2023) highlights that extreme cold events have decreased across much of the U.S., while the frequency, intensity, and duration of extreme heat have increased. In addition, other weather events such as heavy precipitation, drought, floods, wildfires, and hurricanes are becoming more frequent and/or severe, with cascading impacts in all parts of the U.S. (USGCRP, , 2023) and globally (IPCC, 2023). Evaluating their economic impacts requires considering not only local effects but also spillovers across interconnected economic systems (Benzie et al., , 2019; Cashin et al., , 2017).
This study analyses the short-run impact of weather-related disasters, or weather shocks for brevity, on state-level economic activity in the United States. Furthermore, the analysis considers the indirect effects of disasters, namely cross-border spillovers resulting from economic interconnections among states and regions. A monthly Global Vector Autoregressive (GVAR) model (Pesaran et al., , 2004) is employed to capture economic interrelationships between states/regions using bilateral trade flows within the United States. The model incorporates a proxy for weather shocks based on a comprehensive database of emergency declarations triggered by natural disasters. The U.S. represents an interesting case study for assessing the impact of weather-related disasters on the real economy. First, the U.S. territory spans several climate zones, from tropical to polar, and as such it experiences a wide range of extreme weather events, each of which can have a specific impact on the economy. Second, there are a number of reliable indicators that can be taken as proxies for weather shocks, either using geophysical or meteorological datasets, or using damage-based indicators (direct quantification of costs or requests for federal aid by local communities). Third, U.S. data conform to homogeneous standards in terms of how damages are calculated and how aid claims are administered by political authorities. Lastly, the richness of U.S. macroeconomic data makes it possible to study the impact of weather shocks at different levels of geographical disaggregation and to shed light on the transmission channels through which natural disasters affect the real economy.
We find that weather shocks generate direct negative macroeconomic effects at the local level, but also significant negative spillovers to U.S. regions not directly affected. Our results show that weather-related disasters have the potential to act as supply-side shocks, with significant impacts on local economies through disruptions in supply chains and labour markets. The results also indicate that spillovers are stronger when focusing on trade in intermediate goods, thus reinforcing the idea that input linkages across regions and sectors play a key role in amplifying the effects of economic shocks. In addition, we show that geographical disaggregation, parameter heterogeneity and economic linkages between states are key for quantifying spillover effects and the overall impact of weather shocks in the U.S.
@@
We extend the core results along multiple directions to evaluate the robustness and relevance of our findings. We compare our baseline specification with an alternative GVAR model where states are connected by a spatial adjacency matrix, to assess the importance of considering economic interconnections rather than spatial proximity when measuring spillovers. Moreover, since the GVAR model allows for parameter heterogeneity across states, we evaluate the importance of this feature by comparing it with its “homogeneous-parameter counterpart”, namely a spatial panel dynamic model. Next, we consider a country-wide regression for the U.S. to check whether disaggregation at the state level matters at all. Furthermore, we quantify indirect (second-round) effects on local activity for each state, by comparing responses to weather shocks between the GVAR model – that allows for spillovers – and individual state-specific Autoregressive Distributed Lag (ARDL) models where spillovers are implicitly muted. We also address the potential endogeneity of our proxy of weather-related disasters. To this end, we use a two-stage instrumental variable approach that links our indicator to more exogenous, though potentially less precise, proxies for the magnitude of severe weather shocks. These include the number of deaths caused by extreme natural events and a storm indicator that reflects the physical severity of weather phenomena. Lastly, we compare the baseline results with those obtained through the use of two alternative proxies for weather shocks: one based on ex-ante impact estimates generated using the CLIMADA modelling framework (Aznar-Siguan and Bresch, , 2019), and the other based on the Actuaries Climate Index (see, e.g., Kim et al., , 2025).
@@
The paper is related to several streams of literature. First, it adds to the rapidly growing literature on the macroeconomic effects of extreme weather events and climate change (see, e.g., Boldin and Wright, , 2015; Burke et al., , 2015; Colacito et al., , 2019; Dell et al., , 2014; Colombo and Ferrara, , 2024; Billio et al., , 2020; Ahmadi et al., , 2022; Batten, , 2018; Hsiang and Jina, , 2014). In particular, it contributes to the strand of research on cross-border impacts of climate change (Benzie et al., , 2019; Carter et al., , 2021; Feng et al., , 2025). There is ample evidence that the effects of natural disasters propagate through international trade and production networks (Barrot and Sauvagnat, , 2016; Boehm et al., , 2019; Carvalho et al., , 2021; Feng et al., , 2025; Forslid and Sanctuary, , 2023; Kashiwagi et al., , 2021). Our paper is also related to the literature on climate econometrics (see, e.g., Burke et al., , 2015; Hsiang, , 2016; Kahn et al., , 2021). Focusing on the methodology, it is worth pointing out that the GVAR framework has been used to investigate the international spillovers of El Niño weather shocks by Cashin et al., (2017).
@@
Five papers, to the best of our knowledge, are closely related to our work, in that they focus on the effects of climate and weather extremes on the U.S. economy. Mohaddes et al., (2023) and Natoli, (2023) investigate the macroeconomic effects of climate change across the U.S. over the periods 1963–2016 and 1970-2019, respectively. Unlike our analysis, these authors focus on temperature and precipitation anomalies rather than natural disasters, and do not assess spillover effects between states. Kim et al., (2025) investigate the effects of severe weather conditions, proxied by the Actuaries Climate Index, on the aggregate U.S. economy over the period 1963-2019. They find that the effects have become significant only in recent years. Eickmeier et al., (2024) analyze the impact of natural disasters on the aggregate U.S. economy and financial markets using monthly data from 2000 to 2019, and focusing on extreme temperature events, floods, and storms. Relying on local projections and considering the U.S. economy as a whole, they evaluate the impact of adverse weather shocks, measured as the number of natural disasters in the U.S. in a given month, on a large set of macroeconomic and financial variables. Although their analysis does not consider spillover effects, overall they find a significant and persistent negative impact of natural disasters on the real economy, an increase in the price of food and energy, and a reaction of monetary and fiscal policy aimed at containing the negative macroeconomic impacts. Roth Tran and Wilson, (2020) mainly focus on the long-run impact of natural disasters on income using a proxy similar to ours, based on U.S. disaster declarations. Relying on panel-data local projections for annual per capita personal income at county level, they obtain little evidence for a negative effect in the short run, but robust evidence of an increase in the longer run, which is coherent with the “build back better” scenario. Moreover, checking for geographical spillovers, they find scarce evidence that local aid boosts per capita income in the wider region, in the long run. Our paper relies on monthly data and focuses on short-run impacts, thus we complement their analysis based on annual data and concentrating on long-run impacts. Moreover, we study spillover effects based on economic interdependence, while Roth Tran and Wilson, (2020) focus on spatial spillovers.
Overall, our findings reinforce the evidence of a significant economic impact of extreme weather conditions, both at local and aggregate level. Although through different methodologies and different levels of aggregation/disaggregation of the data, our results are in line with other contributions in the literature, like Eickmeier et al., (2024), Colacito et al., (2019), Natoli, (2023), Mohaddes et al., (2023), Kim et al., (2025); this latter finds negative effects mainly for the last part of the sample, which approximately corresponds to our period of investigation. A notable exception is represented by Roth Tran and Wilson, (2020), who instead find slightly negative but statistically insignificant short-run (contemporaneous) responses of real activity indicators to adverse climate shocks.
@@
The rest of the paper is organized as follows. Section 2 presents the data, with a particular focus on our proxy for severe weather shocks; Section 3 discusses the econometric model; Section 4 presents the main results; Section 5 provides further results and robustness checks; Section 6 concludes. An online Appendix with additional results completes the paper.
@@ 2 Data
2.1 Weather-related disasters in the U.S.
We build a novel monthly proxy of weather shocks using the Disaster Declarations Summary dataset maintained by the U.S. Federal Emergency Management Agency (FEMA). FEMA, established in 1979 with an executive order signed by President Carter, is part of the Department of Homeland Security and is responsible for coordinating the federal government’s relief efforts after natural or man-made domestic disasters. All federally declared disasters since 1953 are collected in the Disaster Declarations Summary. For each declaration, the database provides detailed information on the states and counties affected by the disaster, its beginning and end dates, as well as the type of disaster.
@@
Notes: The maps show the count of emergency declarations (total and by type of event). Colors range from lighter to darker as the count of events per state grows. Minimum and maximum values are shown in the color bar. Alaska and Hawaii are omitted to improve visual clarity.
We focus on declarations caused by events that are directly related to weather conditions. In particular, we consider the following main types of events that group the weather-related declarations in FEMA’s database: winter weather (i.e., freezing, ice storms, snow), tropical storms (i.e., hurricanes, typhoons and tornadoes), storms (i.e., coastal and severe storms), fires, floods, and droughts. Most events show the expected seasonal patterns, with storms and fires being the most recurrent types of event. Figure 1 shows the count of declarations by state and by type. Darker colors indicate a higher number of events. If we do not distinguish by event type – upper left panel – we can see that emergency declarations are quite evenly distributed across states. The same holds true for storms and floods. On the contrary, declarations triggered by severe winter weather, tropical storms, fires and drought are geographically concentrated in the northeast, southeast, and west of the U.S., respectively. [^3] Further descriptive statistics on weather events and alternative representations of the maps in Figure 1, either based on the count of declarations weighted by state population density or aggregated over U.S. climate regions, are provided in the online Appendix.
Measuring weather shocks with disaster declarations. We construct our proxy of weather shocks as follows. First, for each state $i$ we create a dummy variable that takes value one if a weather-related disaster was declared in month $t$, and zero otherwise ($emergency_{it}$). Second, we multiply this dummy variable by a weight, given by the ratio between the number of counties affected by the disaster ($hit_{it}$) and the total number of counties (or county equivalents) in state $i$ ($counties_{i}$):
@@
Therefore, our state-level weather shock has unit value only if, in a given month, emergency declarations are issued for all the counties in a state. To calculate an aggregate weather shock for the U.S., we consider the number of counties hit as a share of the total number of counties in the 50 states considered in the analysis ($N_{c}=3142$): [^4] The list of U.S. counties and county equivalents is sourced from: https://en.wikipedia.org/wiki/County.
@@
The weather shock series for the U.S. is displayed in Figure 2, where we highlight the five largest shocks over the period 1990-2019. Hurricane Katrina is by far the largest event, as measured by the share of counties where emergency declarations were issued. Figure 2 also shows that the number of declarations for the U.S. is highly volatile, but there is no visible trend, nor structural breaks in the series. The weighting of events based on the percentage of affected counties in the state is intended to highlight the intensity of disasters rather than the importance of individual counties in the economy of a given state. [^5] In the online Appendix, we consider alternative weighting schemes and proxies of weather shocks. See also Section 5.1.
Figure 2: Weather shocks for the U.S.
@@
Notes: For Hurricane Katrina, emergency declarations were issued in almost all states. In states not directly hit by the hurricane, the declarations go under the header “Hurricane Katrina Evacuation”.
@@
To sum up, the FEMA dataset includes all disasters “of such severity and magnitude that effective response is beyond the capabilities of the state $(\ldots)$ and that supplemental federal assistance is necessary.” [^6] Source: https://www.fema.gov/disaster/how-declared Extreme weather events measured in this way represent “proper” shocks in the parlance of Ramey, (2016): (i) they are exogenous with respect to the current and lagged endogenous variables of the model; (ii) they are uncorrelated with other exogenous shocks (e.g., technological, monetary or fiscal policy shocks); and (iii) they are generally unanticipated.